OpenStudy (flower2468):

1/4, 1/4, 1/4, 1/8, 1/8, 7/12 What is the mean of the data set?

OpenStudy (flower2468):

A. 19/72 B. 19/60 C. 1/20 D. 1 7/12

OpenStudy (flower2468):

I know that it is not D

OpenStudy (anonymous):

so u want mean

OpenStudy (sugardemonhunters):

First, How do you get a Mean?

OpenStudy (flower2468):

you add all the numbers and then divide it by the number of numbers

OpenStudy (dayakar):

\[mean =\frac{ \sum x }{ number of observations }\]

OpenStudy (mathmale):

Yes. You have 6 fractions here. Add these 6 fractions together, after first having obtained the LCD. Then divide this sum by 6.

OpenStudy (anonymous):

\[1/4+1/4+1/4+1/8+1/8+7/12+9/12 /6=19/72\]

OpenStudy (mathmale):

What do you think the LCD is?

OpenStudy (mathmale):

Can you divide 8 into 12 with no remainder?

OpenStudy (flower2468):

no

OpenStudy (mathmale):

Then 12 can't be the LCD. Try again.

OpenStudy (mathmale):

Can you divide 12 into 4 with no remainder?

OpenStudy (mathmale):

You need to find the smallest number divisible (without remainder) by 4, 8 and 12.

OpenStudy (flower2468):

OH I see now

OpenStudy (mathmale):

So, the LCD is: ? Can you divide 2 by 12 with no remainder?

OpenStudy (mathmale):

Again, the 3 separate denoms. are 4, 8 and 12. What is the LCD? Note that 4=2^2, and that 8=2^3, and that 12=4*3=3*2^2

OpenStudy (flower2468):

what is LCD?

OpenStudy (mathmale):

What's the highest power of 2? Answer: 2^3 = 8 What's the highest power of 3? Answer: 3^1=3 What's 8*3?

OpenStudy (flower2468):

8*3=24

OpenStudy (mathmale):

Does 12 divide into 24 with no remainder? Does 8 divide into 24 with no rem? Does 3 div. into 24 with no rem?

OpenStudy (flower2468):

yes

OpenStudy (mathmale):

Great! 24 is your LCD. Convert 1/4 to a fraction with denominator 24, please. 1/4 = ? /24

OpenStudy (flower2468):

1/4= 4*6=24 and 1*6=6 so 6/24

OpenStudy (mathmale):

Cool. Convert 1/8 and 7/12 into fractions with denoms. 24.

OpenStudy (flower2468):

okay 1/8=3/24 and 7/12=14/24

OpenStudy (mathmale):

Cool.

OpenStudy (mathmale):

Now we have to add these 6 fractions together. 1/4 shows up 3 times, 6/24 (same thing) shows up 3 times. What is 3 times 6/24?

OpenStudy (flower2468):

18/24?

OpenStudy (mathmale):

perfect. Hold that 18/24. Next, find 2 times 1/8, which is the same as 2 times 3/24.

OpenStudy (flower2468):

6/24

OpenStudy (mathmale):

Great. Now add together your 3 sums. 18/24 plus 6/24 plus 14/24 = ?

OpenStudy (flower2468):

18+6+14=38 so 38/12

OpenStudy (mathmale):

But our LCD is 24, not 12. What happened there?

OpenStudy (flower2468):

oh right 38/24

OpenStudy (mathmale):

That's it. That's the sum of the given fractions. Any questions?

OpenStudy (flower2468):

no

OpenStudy (flower2468):

wait what about dividing it?

OpenStudy (mathmale):

First compare y our results to the four given answer choices.

OpenStudy (flower2468):

I think that it a

OpenStudy (mathmale):

Is your 38/24 the same as answer A? Don't forget you can reduce 38/24 if necessary.

OpenStudy (flower2468):

38/2=19 and 24/2=12 and well I thought that A sounded more correct then B since it was going to be either A or B

OpenStudy (mathmale):

remember, we were aiming to find the MEAN. You've summed up the 6 numbers. Have you divided this sum by 6?

OpenStudy (flower2468):

no

OpenStudy (mathmale):

Do that, please. That's how we find the mean of these 6 fractions.

OpenStudy (mathmale):

38/24 divided by 6 is what? Note that your answer can and should be reduced.